OpenStudy (anonymous):

Please! Choose the correct simplification of the expression (2x7y)2(y5)3. :)

OpenStudy (anonymous):

Is the expression: \[2x ^{7}y * 2y ^{5} * 3?\]

OpenStudy (anonymous):

its (2x^7)^2(y^5)^3

OpenStudy (anonymous):

to make it easyer for ya 4x^7y^10 4x^9y^10 4x^14y^16 4x^14y^17

OpenStudy (anonymous):

\[(ab)^m=a^mb^m\] \[(a^m)^n=a^{mn}\]

OpenStudy (anonymous):

is it this? 4x^7y^10